# Multiplexed Processing of Quantum Information Across an Ultra-wide Optical Bandwidth

### v3

@@ Abstract
Protocols of quantum information processing are the foundation of quantum technology, allowing to share secrets at a distance for secure communication (quantum key distribution), to teleport quantum states, and to implement quantum computation. While various protocols have already been realized, and even commercialized, the throughput and processing speed of standard protocols is generally low, limited by the narrow electronic bandwidth of the measurement apparatus in the MHz-to-GHz range, which is orders-of-magnitude lower than the optical bandwidth of available quantum optical sources (10-100 THz). We present a general concept and methods to process quantum information in parallel over multiplexed frequency channels using parametric homodyne detection for measurement of all the channels simultaneously, thereby harnessing the optical bandwidth for quantum information in an efficient manner. We exemplify the concept through two basic protocols: Multiplexed Continuous-Variable Quantum Key Distribution (CV-QKD) and multiplexed continuous-variable quantum teleportation. We demonstrate the multiplexed CV-QKD protocol in a proof-of-principle experiment, where we successfully carry out QKD over 23 uncorrelated spectral channels, with capability to detect eavesdropping in any channel. These multiplexed methods (and similar) will enable to carry out quantum processing in parallel over hundreds of channels, potentially increasing the throughput of quantum protocols by orders of magnitude.
@@
In the decades since the conception of quantum information in the 1980s, many practical applications were developed, ranging form secure communications [1, 2, 3] and quantum information transmission protocols [4], through quantum sensing schemes [5, 6, 7, 8] to quantum computation [9, 10, 11, 12]. All these applications rely on some uniquely quantum properties, such as entanglement, squeezing, etc., to encode, process and decode the desired quantum information. Examples include various degrees of freedom of matter, such as the energy levels of trapped ions [11] and neutral atoms [12] or the phase / charge of Josephson junctions [13, 14, 15]; and of light, either in the discrete photon-basis (e.g. polarization [1, 2], spatial mode [16], frequency [17] and time [18]) or in the continuous quadrature-basis (squeezing [6, 19]).
@@
In many applications, the quadratures of the optical electric field are the backbone of optical quantum processing. Classically, the quadratures $x,y$ are the cosine and sine components of the optical field at frequency $\omega$ $E(t)=x\cos\left(\omega t\right)+y\sin\left(\omega t\right)=\left|a\right|\cos\left(\omega t+\varphi\right)$, where $a=\left|a\right|e^{i\varphi}$ is the complex field amplitude. The quadratures are the real $x=a+a^{*}$ and imaginary $y=i\left(a-a^{*}\right)$ components of the complex amplitude. In quantum optics, the optical quadratures ($x=\left(a+a^{\dagger}\right)/2,y=i\left(a-a^{\dagger}\right)/2$) are non-commuting observables $[x,y]=i/2$ that maintain the canonical quantum uncertainty $\Delta x\Delta y\geq 1$, analogous to the canonical position and momentum of quantum mechanics. One can therefore encode, store, process and decode quantum information on the optical quadratures.
@@
The focus of this paper is to harness the bandwidth of ultra-broadband optical sources to drastically enhance the rate of quantum optical processing. Most generally, quantum information processing can be broken into three primary stages: generation of the quantum state, manipulation of the state, and its measurement. While the speed of each stage can be limited by different factors, the primary bottleneck in quantum optical protocols is the measurement, where the relatively slow response of photo-detectors limits the processing rates at several orders-of-magnitudes below the optical bandwidth of available sources, even with the fastest available detectors. In particular, sources of broadband squeezed light with 10-100THz of bandwidth (up to an optical octave) are readily available [20, 21], as well as methods of broadband manipulation using pulse shaping in the spectral domain [22, 23]. In contrast, the bandwidth of traditional measurement techniques was always limited by the narrowband electronic response of optical detectors, in the MHz-to-GHz range.
@@
Luckily, this electronic bandwidth limit was recently overcome with the conception of optical parametric homodyne [24], which enables to measure an optical quadrature of interest across a wide, practically unlimited optical spectrum, opening the way to much faster quantum processing. We present a general approach for parallel processing of quantum information, encoded across the entire optical spectrum of the quadratures of broadband two mode squeezed light. We highlight a set of tools (see figure 1) to simultaneously generate, manipulate and measure quantum information over multiple frequency channels, up to $10^{3}-10^{4}$ channels in realistic configurations, limited only by the available optical bandwidth. Here, these tools helped us to develop a multiplexed quantum teleportation protocol, which can teleport multiple quantum states simultaneously, as well as a multiplexed QKD protocol (BB84-like), which we demonstrated experimentally over 23 spectral channels in parallel. We note however that the presented toolkit is useful far beyond those two examples, and can be used to form multiplexed variations of any existing quantum protocol, thereby enhancing the processing throughput by orders of magnitude.
@@
Figure 1: The tool-set for multiplexed processing with broadband squeezed light, decomposed to generation, manipulation and detection of quantum states in a multiplexed manner. Later we employ this tool-set to assemble two quantum protocols (multiplexed QKD and quantum teleportation), as two major examples. For state generation we use a broadband optical parametric amplifier (OPA), which generates a broad spectrum of two-mode squeezed light from a narrowband pump. For state manipulation we use spectrally-resolved phase shifters and beam splitters, which allow to control the quadratures of the quantum states separately for each channel. Finally, for detection we employ parametric homodyne using another OPA followed by a spectrometer, where the chosen quadrature is amplified and simultaneously measured across the spectrum.
.1 Multiplexed QKD
As a first example for a multiplexed protocol, let us present (and later demonstrate experimentally) a simple protocol of multiplexed QKD. We note up front that this protocol is not intended as an immediate practical implementation of ultrafast QKD, but rather as an illustration of the viability of frequency-multiplexed quantum processing. Our scheme forms a continuous-variable analog of the BB84 protocol using an unseeded SU(1,1) interferometer. Specifically, both Alice and Bob have an unseeded OPA that generates broadband SPDC and a broadband phase modulation device that consists of a Fourier-domain spectral shaper (see figure 2 and caption for details). Since an unseeded SU(1,1) interferometer generates a wide spectrum of signal-idler pairs, the different frequencies within this spectrum can be used as separate QKD channels.
When a pump laser passes through two OPAs in series, the SPDC generation in the 2nd OPA can interfere with the SPDC generation in the 1st OPA, depending on the phase of the signal-idler pair relative to the pump. This leads us to the following 4-steps protocol:
@@
    To encodes her information, Alice modulates the phase of each signal-idler channel in one of two mutually unbiased bases (chosen at random): Basis 1 uses $\phi=0$ (constructive interference) for logic ’1’ and $\phi=\pi$ (destructive) for logic ’0’, whereas basis 2 employs $\phi=\pm\frac{\pi}{2}$. After the spectral modulation of all the channels (in parallel), Alice sends the phase modulated spectrum to Bob.
@@
    To detect the information, Bob randomly chooses a measurement basis (for each channel separately) using his spectral shaper - by setting the phase to 00 for basis 1 or to $\frac{\pi}{2}$ for basis 2, and passes the light again through his OPA, where the SU(1,1) interference occurs. Bob measures the spectrally resolved light intensity with a spectrometer, which reflects the number of photons in each channel at the output of the complete SU(1,1) interferometer. If Bob sets the correct basis for a channel, the interference of that channel at the output will be either fully constructive (high probability for photo-detection) or destructive (low probability) and Bob will be able to detect Alice’s phase. However, if Bob sets the phase to the wrong basis, his interference will be intermediate, preventing Bob from deducing the information.
@@
    Finally, to detect a possible eavesdropper, Alice and Bob compare a fraction of their data, searching for errors that Eve’s measurements would have introduced (just like any other QKD protocol).
Figure 2: The multiplexed QKD protocol. To encode the data Alice employs a source of broadband, weakly squeezed light (entangled photon pairs) generated using an OPA with broadband phase-matching. Alice modulates the spectral phase of the squeezed light using a Fourier-domain spectral shaper that consists of a spectral separator, such as grating or prism and a spatial light modulator (SLM) to vary the phase of each spectral channel independently. The modulated light is then sent to Bob via an optical fiber. To decode Alice’s information, Bob chooses a measurement basis for each channel by modulating the spectral phase of the squeezed light with his spectral shaper (identical to Alice’s) and passes the light through a second OPA (completing the SU(1,1) interferometer). Finally, Bob measures the spectral intensity with a spectrometer trying to detect single photons in each channel.
The security of each channel within this scheme can be analyzed similar to the standard analysis of the BB84 protocol, as summarized hereon (the complete derivation of the security is given in the methods - section I.1). Assuming a weak parametric gain in the OPAs, we can employ the perturbative quantum propagator through the nonlinear crystal [25], as $U(t)=e^{iHt}\approx 1+iHt=1+ig_{\omega}a^{\dagger}_{\omega}a^{\dagger}_{-\omega}$, where $a_{\pm\omega}$ represent the field operators of the signal-idler mode pair at $\omega_{p}/2\pm\omega$ and $g_{\omega}$ represent the parametric gain of that pair, which includes the interaction time, $t$ (proportional to the crystal length). Assuming the vacuum state as the input to Alice’s OPA, the output state after Bob’s OPA is
@@
where $\phi_{\omega}=\phi_{A}+\phi_{B}$ is the total phase that Alice and Bob apply to the signal-idler pair (relative to the pump field) during steps 1 and 2 of the protocol. The phase $\phi_{A(B)}$ indicates the basis of encoding (measurement) that Alice (Bob) employ for each bit of the channel. The average number of photon that Bob will measure at a specific channel is
$N_{\omega}=|\left<1_{\omega}|\psi\right>_{2}|^{2}=|g_{\omega}|^{2}(2+2\cos(\phi_{\omega}))$ (2)
Notice that when Alice and Bob use different bases, $\phi_{\omega}=\phi_{A}+\phi_{B}=\frac{\pi}{2}/\frac{3\pi}{2}$ and the average number of photons is simply $2|g_{\omega}|^{2}$, independent of the phase. However, when the bases match, $\phi_{\omega}=\phi_{A}+\phi_{B}=\pi/0$ the average number of photon equals $0/4|g_{\omega}|^{2}$ respectively, so Bob’s measurement can decode Alice’s information.
When Eve tries to attack the communication the situation changes noticeably. For example, if Eve ”steals” some of the light using a beam splitter with transmission $T$, then the number of photons after Bob’s OPA becomes (see derivation in the methods, section I.1.1)
@@
$V(T)\equiv\frac{I_{max}-I_{min}}{I_{max}+I_{min}}=\frac{2T}{1+T}.$ (4)
@@
The contrast is a witness for eavesdropping since Eve must steal some of the photons, i.e. reduce the transmission ($T$) which will lower the contrast. Notice that Eve’s ability to extract information relies on a similar interference contrast (in her own measurements), $V_{Eve}=V(R=1\!-\!T)=\frac{2(1-T)}{2-T}$. Thus, to obtain a sufficient contrast in her measurements, Eve must induce a sufficiently high loss in her beam splitter, which Alice and Bob can later identify.
.2 Multiplexed Quantum Teleportation

### v4

@@ Abstract
Protocols for processing of quantum information are the foundation of quantum technology, enabling to share secrets at a distance, teleport quantum states, and to implement quantum computation. While many protocols were realized, and even commercialized, the throughput and processing speed of current protocols is limited by the narrow electronic bandwidth of standard measurement devices (typically in the MHz-to-GHz range), which is orders-of-magnitude lower than the optical bandwidth of available quantum optical sources (10-100 THz), indicating that the bandwidth resource is dramatically underutilized in current quantum optical technology. We present a general concept of frequency multiplexed quantum channels and a set of methods to process quantum information efficiently across the available optical bandwidth. Using a broadband source of squeezed light, spectral manipulation methods and parametric homodyne detection, we are able to generate, process and measure all the channels in parallel, thereby harnessing the optical bandwidth for quantum information in an efficient manner. We exemplify the concept through two basic protocols: Multiplexed Continuous-Variable Quantum Key Distribution (CV-QKD) and multiplexed continuous-variable quantum teleportation. The multiplexed QKD protocol is demonstrated in a proof-of-principle experiment, where we successfully carry out QKD over 23 uncorrelated spectral channels, with capability to detect eavesdropping in any channel. These multiplexed methods (and similar) will enable to carry out quantum processing in parallel over hundreds of channels, potentially increasing the throughput of quantum protocols by orders of magnitude.
journal: opticajournalarticletype: Research Article
1 Introduction
@@
In the decades since the conception of quantum information in the 1980s, many practical applications were developed, ranging form secure communications [1, 2, 3] and transmission of quantum information [4], through quantum sensing schemes [5, 6, 7, 8] to quantum computation [9, 10, 11, 12]. All these applications rely on some uniquely quantum property, such as entanglement, squeezing, etc., to encode, process and decode the desired quantum information. Practically any degree of freedom of light can carry quantum information, in either discrete photons or in continuous variables, such as polarization [1, 2], spatial mode [13], field quadratures [6, 14], frequency [15] and time [16].
@@
Specifically, the quadratures of the optical electric field are the backbone of many protocols of optical quantum processing. Classically, the quadratures $x,y$ are the cosine and sine components of the optical field at frequency $\omega$: $E(t)=x\cos\left(\omega t\right)+y\sin\left(\omega t\right)=\left|a\right|\cos\left(\omega t+\varphi\right)$, where $a=\left|a\right|e^{i\varphi}$ is the complex field amplitude. The quadratures are the real $x=a+a^{*}$ and imaginary $y=i\left(a-a^{*}\right)$ components of the complex amplitude. Quantum mechanically, the optical quadratures are analogous to the canonical position and momentum, defined as $x=\left(a+a^{\dagger}\right)/2,y=i\left(a-a^{\dagger}\right)/2$ with the canonical commutation relation $[x,y]=i/2$ and uncertainty relation $\Delta x\Delta y\geq 1$. One can therefore encode, store, process and decode quantum information on the optical quadratures. Specifically, squeezed states of light, where the fluctuations of one quadrature are squeezed below the vacuum level, while the orthogonal quadrature is stretched, can be utilized for quantum processing and represent quantum-bits (qubits) at both low and high squeezing levels, as we show hereon.
@@
The focus of this paper is to harness the bandwidth of ultra-broadband sources of squeezed light to drastically enhance the rate of quantum optical processing. Most generally, quantum information processing can be broken into three primary stages: generation of the quantum state, manipulation of the state, and its measurement. While the speed of each stage can be limited by different factors, the primary bottleneck in quantum optical protocols is the measurement, where the relatively slow response of photo-detectors limits the processing rates at several orders-of-magnitudes below the optical bandwidth of available sources, even with the fastest available detectors. In particular, sources of broadband squeezed light with 10-100THz of bandwidth (up to an optical octave) are readily available [17, 18], as well as methods of broadband manipulation using pulse shaping in the spectral domain [19, 20]. In contrast, the bandwidth of traditional measurement techniques was always limited by the narrowband electronic response of optical detectors, in the MHz-to-GHz range.
@@
Luckily, this electronic bandwidth limit was recently overcome with multimode SU1,1 interferometry and with parametric homodyne detection [21], which enable to measure an optical quadrature of interest across a wide, practically unlimited optical spectrum, opening a path to much faster quantum processing. We present a general approach for parallel processing of quantum information, encoded across the entire optical spectrum of the quadratures of broadband two mode squeezed light. We highlight a set of tools, illustrated in figure 1, to simultaneously generate, manipulate and measure quantum information over multiple frequency channels, up to $10^{3}-10^{4}$ channels in realistic configurations, limited only by the available optical bandwidth.
@@
Figure 1: The tool-set for multiplexed processing with broadband squeezed light, decomposed to generation, manipulation and detection of quantum states in a multiplexed manner. State generation employs a broadband optical parametric amplifier (OPA), which generates a broad spectrum of two-mode squeezed light from a narrowband pump, either by spontaneous parametric down-conversion (SPDC) or four-waves mixing (FWM). State manipulation is performed by spectrally-resolved phase shifters and by beam splitters, which allow to control the quadratures of the quantum states separately for each channel. Finally, the detection employs parametric homodyne in an additional OPA, where the chosen quadrature is amplified, followed by a spectrometer for simultaneous spectrally resolved measurement of that quadrature. Later we employ this tool-set to assemble two quantum protocols (multiplexed QKD and quantum teleportation), as two major examples.
Here, the tools of figure 1 helped us to develop a multiplexed version for two of the most important quantum processing protocols:
@@
    Multiplexed quantum teleportation; Teleportation is a critical stepping stone of quantum computation in particular and of quantum technologies in general. For example, in measurement-based quantum computation, which is the chosen scheme for leading companies, such as PsiQuantum [22] and Xanadu [23], quantum teleportation is key for implementing all 2-qubit gates [24, 25]. Thus, multiplexed quantum teleportation, as we present hereon can significantly scale-up the number of qubit-gates and speed up the processing accordingly.
@@
    Multiplexed QKD (BB84-like); QKD provides safe key sharing between two distant parties, whose security is physically-proven. Although many QKD protocols were already demonstrated, the communication speed in all of them is limited by the detection bandwidth of a single channel, which is bound by electronic performance. Our QKD protocol can bridge the orders of magnitude gap between the wide bandwidth of light sources and the narrow bandwidth of detectors. To highlight the capabilities of our set of tools, we demonstrate our multiplexed QKD in a proof-of-principle experiment, which successfully processed information over 23 parallel spectral channels.
And yet, those two protocols serve primarily to exemplify the general utility of the presented toolkit - to multiplex quantum protocols with entangled photons / squeezed light. In general, we expect our set of tools to be useful also for many other quantum processing technologies thereby enhancing their processing throughput by orders of magnitude.
2 Multiplexed QKD
As a first example for a multiplexed protocol, let us present (and later demonstrate experimentally) the simple protocol of multiplexed QKD with entangled bi-photons, illustrated in figure 2. Our scheme forms a continuous-variable analog of the BB84 protocol [1, 26, 27] using an unseeded SU(1,1) interferometer [28], where two OPAs are set in series and pumped by a single pump. The phase-sensitive OPAs amplify / attenuate one quadrature of light with no added noise [21, 28], indicating that the generation of bi-photons in the two OPAs can fully interfere for any gain settings, depending on the phase of the signal-idler pairs relative to the pump in the second OPA. The pump laser therefore serves as the local oscillator to decode the phase of each signal-idler pair.
In our protocol, Alice holds the first OPA of the interferometer and Bob holds the second. Alice’s OPA is unseeded (vacuum input) and operates at low gain, generating randomly time-energy entangled bi-photons through either spontaneous parametric down-conversion (SPDC) or spontaneous four-waves mixing (SFWM). Alice encodes information on the broad spectrum of bi-photons from her OPA using a Fourier-domain spectral shaper that modulates the spectral phase of the light using a spatial phase modulator. Bob can then complete the interferometer with his OPA, thereby allowing him to measure the encoded phase (see figure2 and caption for details). Since an unseeded OPA with a broad phase-matching generates a wide spectrum of signal-idler pairs, the different frequencies within this spectrum can be used as separate QKD channels.
This leads us to the following 4-steps protocol with the configuration of figure 2:
  1. 1.
    To encode information, Alice modulates the phase of each signal-idler channel in one of two mutually unbiased bases (chosen at random): Basis 1 uses $\phi=0$ (constructive interference) for logic ’1’ and $\phi=\pi$ (destructive) for logic ’0’, whereas basis 2 employs $\phi=\pm\frac{\pi}{2}$. After the modulation of the spectral phase of all the channels in parallel (e.g. using a Fourier-domain pulse shaper) Alice sends the phase modulated spectrum to Bob together with a phase reference (e.g. her pump laser).
  2. 2.
    To detect the information, Bob randomly chooses a measurement basis (for each channel separately) using another spectral phase modulator - by setting the phase to 00 for basis 1 or to $\frac{\pi}{2}$ for basis 2 (relative to the phase reference he received from Alice). The, Bob passes the light again through his OPA, where the SU(1,1) interference occurs. Bob measures the spectrally resolved light intensity with a spectrometer, which reflects the number of photons in each channel at the output of the complete SU(1,1) interferometer. If Bob sets the correct basis for a channel, the interference of that channel at the output will be either fully constructive (high probability for photo-detection) or destructive (low probability) and Bob will be able to detect Alice’s phase. However, if Bob sets the phase to the wrong basis, his interference will be intermediate for both ’0’ or ’1’, preventing Bob from deducing the encoded information.
@@ 2 Multiplexed QKD
    Finally, to detect a possible eavesdropper, Alice and Bob compare a fraction of their data, searching for errors that Eve’s measurements could have introduced (just like any other QKD protocol).
This 4-steps protocol formulates the physical layer of a multiplexed QKD communication, i.e. forms a baseline of spectral channels on which QKD can be performed by various protocols of data-exchange. Specifically, a higher protocol should define logical components that are critical for a practical QKD implementation, such as clock synchronization and error-mitigation. For example, the ambiguity between the destructive interference event (bit ’0’) and the no-photon event (communication fail) can be mitigated if Bob divides his integration time into two time-windows, and flips the measurement phase in the middle by $\pi$ (similar to differential signaling in classical communication). Thus, a successful detection event for Bob must include one photon in exactly one of the time-halves (the logical ’0’ or ’1’ is dictated by their order). This allows Bob to identify errors and discard no-photon events. As there are many possible ways to implement such higher-level protocols, we will leave their optimization to future research.
Figure 2: The multiplexed QKD protocol. To encode the data Alice employs a source of broadband entangled photon pairs (two-mode, weakly squeezed light [29]), generated using an OPA with broadband phase-matching. Alice modulates the spectral phase of the squeezed light using a Fourier-domain spectral shaper that consists of a dispersive element, such as grating or prism to spatially separate the frequencies and a spatial light modulator (SLM) to vary the phase of each spectral channel independently. The modulated light is then sent to Bob via an optical fiber along with the pump that serves as a local oscillator for all the frequencies. To decode Alice’s information, Bob chooses a measurement basis for each channel by modulating the spectral phase of the squeezed light with his spectral shaper (identical to Alice’s) and passes the light through a second OPA (completing the SU(1,1) interferometer). Finally, Bob measures the spectral intensity with a spectrometer trying to detect single photons in each channel.
The security of each channel within our scheme can be analyzed similar to the standard analysis of the BB84 protocol, as summarized hereon (the complete derivation of the security is given in supplement 1). Assuming a weak parametric gain in the OPAs, we can employ the perturbative quantum propagator through the nonlinear crystal [30], as $\hat{U}(t)=e^{i\hat{H}t}\approx 1+i\hat{H}t=1+ig_{\omega}\hat{a}^{\dagger}_{\omega}\hat{a}^{\dagger}_{-\omega}$, where $\hat{a}_{\pm\omega}$ represent the field operators of the signal-idler mode pair at $\omega_{p}/2\pm\omega$ and $g_{\omega}$ is the complex gain of that pair , which includes the interaction time $t$ (proportional to the crystal length) and the relative phase between the pump and the pair. Assuming the vacuum state as the input to Alice’s OPA, the output state after Bob’s OPA is
@@
where $\phi_{\omega}=\phi_{A}+\phi_{B}$ is the total phase that Alice and Bob apply to the signal-idler pair (relative to the pump field) during steps 1 and 2 of the protocol. The phase $\phi_{A(B)}$ indicates the basis of encoding (measurement) that Alice (Bob) employ for each bit of the channel. The average number of photons that Bob will measure at a specific channel is
$N_{\omega}=|\left<1_{\omega}|\psi\right>_{2}|^{2}=|g_{\omega}|^{2}(2+2\cos(\phi_{\omega})),$ (2)
where for simplicity we assumed that Bob’s parametric gain is identical to Alice (generalization to nonidentical gain is simple, see [31]). Notice that when Alice and Bob use different bases, $\phi_{\omega}=\phi_{A}+\phi_{B}=\frac{\pi}{2}/\frac{3\pi}{2}$ and the average number of photons is simply $2|g_{\omega}|^{2}$, independent of the phase. However, when the bases match, $\phi_{\omega}=\phi_{A}+\phi_{B}=\pi/0$ and the average number of photon equals 00 or $4|g_{\omega}|^{2}$ respectively, so Bob’s measurement can decode Alice’s information.
Our analysis so far assumed the parametric gain to be low and neglected the generation of multiple photon pairs per bit. Generally speaking, multiple pairs are not desirable, as they open the door to splitting attacks, but note that multiple pairs do not hamper the communication protocol or introduce errors. Indeed, the phase of each frequency (or photon) is random and varies with time, but the phase-sum of the signal-idler pair is deterministically fixed by the pump (and propagation). Since our protocol relies only on the phase-sum, no phase ambiguity will arise, even in the case of multiple photons per bit.
When Eve tries to attack the communication the situation changes noticeably. In the supplementary material we analyze some of the possible realistic attacks that Eve may try. However, in order to illustrate Eve’s detectability, let us consider the simplest possible example - the steal attack. If Eve "steals" some of the light using a beam splitter with transmission $T$, then the number of photons after Bob’s OPA becomes (see derivation in supplement 1)
@@
$V(T)\equiv\frac{I_{max}-I_{min}}{I_{max}+I_{min}}=\frac{2T}{1+T},$ (4)
@@
where again the parametric gains of Alice and Bob are assumed equal (for non equal gains the visibility is $V(T)=\frac{2g_{A}g_{B}T}{g^{2}_{B}+Tg^{2}_{A}}$ [31] ). The contrast is a witness for eavesdropping since Eve must steal some of the photons, i.e. reduce the transmission ($T$) which will lower the contrast. Notice that Eve’s ability to extract information relies on a similar interference contrast (in her own measurements), $V_{Eve}=V(R=1\!-\!T)=\frac{2(1-T)}{2-T}$. Thus, to obtain a sufficient contrast in her measurements, Eve must induce a sufficiently high loss in her beam splitter, which Alice and Bob can later identify. Although the steal attack above is simplistic, and may even be detected by measuring the average intensity, the contrast proves to be a very effective discriminator for Eve’s presence, even with much more realistic and sophisticated attacks (some of which are analyzed the supplementary material).
It is important to understand that our protocol encapsulates a set of single QKD communication channels that are otherwise independent, which indicates that standard QKD constraints apply at the single channel level and not at the entire protocol level. As an example, our spectrally multiplexed QKD allows for a much higher total photon flux (and much higher total data rate) than the equivalent single channel QKD. Specifically, for QKD to be secure, each transmitted bit should contain less than one photon per detection time, but since the spectral channels carry the QKD exchange independently, this low-flux requirement holds per channel. Thus, the total photon generation rate (parametric gain) should be adjusted to produce less than one photon on average is sent per channel bandwidth per detection time. The total flux of photons across all the channels can be much more than one photon.
Clearly our protocol requires Alice and Bob to share a classical phase reference and to be able to modulate the spectral phase relative to it. While figure 2 shows one possible implementation to meet these requirements (by sending the pump along with the SPDC as a common phase reference and using a Fourier domain pulse shaper as a spectral phase modulator), other implementations are possible that may prove technically advantageous in specific configurations. We will briefly consider some of those possibilities in the discussion section later on.
3 Multiplexed Quantum Teleportation